OpenStudy (anonymous):

The length of a rectangle is 8 cm greater than its width. Find the dimensions of the rectangle is its area is 105 cm^2. Write an equations and solve by factoring.

OpenStudy (l):

width = W length = L = W+8 Area = W . L = W . (W+8) = 105 W^2 + 8W = 105 W^2 + 8W - 105 = 0 W^2 + 8W - 105 = (W+15) . (W-7) = 0 -> (W+15) = 0 or (W-7) = 0 -> W = -15 or W = 7 -15 is no a metric value, because it is negative So the width is 7cm W = 7cm L = 7+8 = 15cm *